A coin bank has $37 dollars consisting of quarters and dimes. There are 10 times as many dimes more than 2 times as many quarters. how many quarters and dimes are there
Answer by rapaljer(4671)

Let x = number of quarters
2x+10 = number of dimes

The equation is based upon the values of each coin times the number of each coin. You can do it in either dollars or cents. Your teacher may have a preference, but to me, working with dollars requires the use of decimals, whereas the use of cents comes out in whole numbers. I'll do it in "cents". Doesn't that make "cents" to you?? Of course the quarters are each worth 25 and the dimes are each worth 10 cents. Be sure to change the $37 to cents also.

25(x) + 10(2x+10) = 3700
25x + 20x + 100 = 3700
45x +100 = 3700
45x=3600
x=3600/45=80 Quarters
2x+10 = 2(80)+10= 170 Dimes

Check by calculating the values of the coins:
Quarters = 80*.25 = $20
Dimes = 170*.1= $17
Total = $37 It checks!!

For additional help with coin problems or any kind of word problems, please see my website by clicking on my tutor name "Rapaljer" anywhere in algebra.com. Look for "Basic, Intermediate, and College Algebra: One Step at a Time." Select "Basic Algebra", and look for the appropriate topic in "Chapter 1". This is a user friendly explanation with examples and exercises integrated into it. My students ALL tell me it's easier to understand than the traditional textbooks! See also my "MATH IN LIVING COLOR" pages where a lot of these exercises are solved in "Living Color."

R^2